# Multi-Knowledge-oriented Nighttime Haze Imaging Enhancer for Vision-driven Intelligent Systems

### v2

@@ 5.1.4 Experiment Platform
Figure 7: isual comparisons of detail enhancement.
5.2 Synthetic Degradation Analysis
5.2.1 Dehazing
@@
Table 1 presents the comparative performance of MKoIE against seven ID methods on the RESIDE and CDD datasets. MKoIE achieved the highest average PSNR and SSIM scores, and the second best NIQE score, demonstrating its superiority in quantitative evaluation. Furthermore, MKoIE exhibits substantial advantages in qualitative comparisons. Fig. 5 illustrates the results of various methods on synthetic hazy images. The traditional DCP algorithm enhances contrast and removes some haze but often introduces color distortions and over-dehazes. Deep learning-based methods, such as FFANet and AoSRNet, improve haze removal and detail recovery but exhibit limitations when handling distant objects in natural scenes. DehazeFormer frequently produces prominent artifacts, particularly in sky regions, while CEEF, though capable of partial restoration, suffers from local distortions and loss of detail. Notably, MKoIE, despite being trained on diverse low-visibility images, demonstrates competitive performance comparable to state-of-the-art daytime ID methods, effectively addressing the challenges posed by synthetic and real-world hazy scenarios.
5.2.2 Low-light Enhancement
@@
Table 2 and Fig. 5 illustrate that MKoIE significantly outperforms other methods in LLIE. Specifically, MKoIE demonstrates superior performance in image clarity and detail restoration, producing images with higher fidelity in terms of structural integrity while effectively reducing blurring and artifacts. LIME enhances visibility with noticeable improvements in color contrast and haze removal but introduces slight color distortions. SDD, while improving saturation in building areas and achieving a warmer appearance, suffers from blur and detail loss, particularly at the edges of buildings and within crowded regions. Both SMNet and EFINet exhibit over-sharpening effects, causing certain image details to appear excessively pronounced. In contrast, MKoIE generates images that closely resemble real-world scenes, with textures and colors more consistently preserved, thereby achieving a balance between enhancement and realism.
5.2.3 Nighttime Haze Enhancement
@@
The evaluation of MKoIE was conducted using the RESIDE and CDD datasets, consistent with those in daytime ID and LLIE tasks. As shown in Fig. 5, while IENHC and ECNHI partially remove haze, residual haze remains visible, particularly in the background regions. In contrast, MKoIE cexhibits exceptional performance in color restoration, producing images with a natural tone devoid of oversaturation or color distortion, thereby enhancing both visual comfort and scene realism. In addition, MKoIE demonstrates superior detail preservation, with ground textures in the foreground and building outlines in the background clearly discernible. Quantitative evaluations in Table 3 further substantiates MKoIE’s superiority, as it achieves the best results in all metrics, underscoring its outstanding recovery performance under low-light and hazy conditions. Figure 6’s histogram shows that our method achieves a balanced intensity distribution, along with superior brightness, color, and contrast enhancements. MKoIE produces a broad, well-spread histogram, indicating an optimal use of the intensity range. This corresponds to better brightness, more natural color distribution, and higher contrast compared to the other methods. Similarly, we conducted a detail enhancement comparison experiment, as shown in Figure 7. The original image contains significant haze, especially around the object. IENHC yields a moderate clarity improvement: text becomes more legible, but background details remain largely unrecovered. MRP and OFSD both reduce haze and enhance image contrast and color saturation, though their results exhibit a slight remaining color cast. CEEF achieves only modest haze removal with a mild increase in contrast. In contrast, MKoIE shows pronounced improvements in overall color, contrast, and brightness. It also excels at restoring fine details, as evidenced by the clear textures visible in the red-boxed zoomed region in Figure 7. These visual enhancements correspond with MKoIE’s top quantitative metrics in Table 3, confirming the method’s superior performance across both objective and subjective evaluations.
5.3 Real-world Degradation Analysis
@@
Figure 8: Zooming-in is recommended for a better visualization and comparisons
@@
Table 4: Ablation Studies on the TNL in ID, LLIE and NHIE tasks.
@@
Table 5: Ablation Studies on the MRFE in ID, LLIE and NHIE tasks.
@@
Table 6: Ablation Studies on the Different Loss Function in ID, LLIE and NHIE tasks.
@@ 5.4 Ablation Study
5.4.1 Ablation Studies on TNL
To validate the effectiveness of the proposed TNL modules, we assign different image restoration tasks to NHIE sub-node to train another version of MKoIE, as shown in Fig. 9, only a single sub-network was used to complete the three tasks. As shown in Table 4, compared to the full version of MKoIE, the performance of ID, LLIE, and NHIE degrades when the TNL module is specified independently. This highlights the effectiveness of the proposed shared and task-specific feature extraction methods. By leveraging its distinct architecture, the TNL module is designed to enhance feature extraction and thus improve overall task performance. It integrates the fundamental features required for ID, LLIE, and NHIE, without TNL module, MKoIE exhibits poor performance in the three tasks.
5.4.2 Ablation Studies on Network Modules
To assess the significance of the MRFE module in MKoIE, we performed an ablation study by completely removing it. As shown in Table 5, its removal consistently led to declines in PSNR and SSIM across all tasks. This demonstrates that the absence of the MRFE module limits the network’s ability to extract essential features, resulting in reduced restoration performance.
5.4.3 Ablation Studies on Loss Function
Each component of the hybrid loss function show in subsection 4.6. To demonstrate the effectiveness of hybrid loss function, we train two versions: one using only the reconstruction loss and the other using only the perceptual loss. From Table 6, the declines in both PSNR and SSIM indicate that hybrid loss function guide the network toward improved restoration performance. Reconstruction and perceptual loss enhances restoration performance by balancing numerical stability, outlier resilience, perceptual consistency, and semantic preservation, achieving optimal results.
5.5 Improving SOD Tasks with MKoIE
Table 7: The comparison for SOD task between original images and images enhanced by the MKoIE includes MAE, max F-measure($F_{\beta}^{\text{max}}$), max E-measure($E_{\phi}^{\text{max}}$)and S-measure($S_{\alpha}$).
@@ 5.5 Improving SOD Tasks with MKoIE
We utilized U2Net [47] and show its performance on CDD dataset. Notably, we directly applied the pretrained U2Net model without further training, achieving a marked enhancement in SOD through MKoIE preprocessing. As shown in Fig. 10 and Table 7, the U2Net model achieved significant improvements in MKoIE-processed images, enabling more accurate detection and segmentation of salient objects. This result validates the effectiveness of MKoIE, demonstrating its practical application value in traffic scenarios under complex weather conditions. Specifically, in autonomous driving and traffic monitoring systems, it enhances the detection and segmentation performance of critical targets such as pedestrians and vehicles, thereby further improving traffic safety.
5.6 Complexity and Running Time Comparison
Table 8: We selected the top four models based on the quantitative comparison of the NHIE task and present the corresponding average runtime (in seconds).
@@ 5.6 Complexity and Running Time Comparison
MKoIE adopts an efficient structural design, which can significantly improve the image recovery speed in low light and haze environments. We randomly selected 50 images from the test set and ran all models on the same hardware environment. As shown in Table 8, the average processing time of our model is about 1.68 seconds. The results show that MKoIE has advantages in both efficiency and effectiveness, demonstrating its potential value in practical applications.
6 Conclusions
This paper proposes a multi-knowledge-oriented imaging enhancer (MKoIE) to address salient object detection challenges in vision measurement systems under adverse imaging conditions. MKoIE effectively handles three types of degradation (i.e., daytime haze, low light, and nighttime haze) through a task-oriented node learning mechanism and self-attention module, while achieving efficient multi-scale feature extraction through a multi-receptive field enhancement module. Experimental results demonstrate that the proposed MKoIE achieves superior performance across various weather and imaging conditions, thus providing an effective solution for image enhancement and salient object detection in challenging environments.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
The work described in this paper is supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (Grant Number. PolyU 15201722), the Key Research and Development Project of Sichuan (Grant Number. 2024YFG0008), and the Major Program of National Natural Science Foundation of China (Grant Number. T2293771).

### v3

@@ V-A4 Experiment Platform
Figure 7: Visual comparisons on synthetic degradation scene. Zooming-in is recommended for a better visualization and comparisons.
V-B Synthetic Degradation Analysis
V-B1 Dehazing
@@
Table I presents the comparative performance of MKoIE against seven ID methods on the RESIDE and CDD datasets. MKoIE achieved the highest average PSNR and SSIM scores, and the second best NIQE score, demonstrating its superiority in quantitative evaluation. Furthermore, MKoIE exhibits substantial advantages in qualitative comparisons. Fig. 5 illustrates the results of various methods on synthetic hazy images. The traditional DCP algorithm enhances contrast and removes some haze but often introduces color distortions and over-dehazes. Deep learning-based methods, such as FFANet and AoSRNet, improve haze removal and detail recovery but exhibit limitations when handling distant objects in natural scenes. DehazeFormer frequently produces prominent artifacts, particularly in sky regions, while CEEF, though capable of partial restoration, suffers from local distortions and loss of detail. Notably, MKoIE, despite being trained on diverse low-visibility images, demonstrates competitive performance comparable to state-of-the-art daytime ID methods, effectively addressing the challenges posed by synthetic and real-world hazy scenarios.
V-B2 Low-light Enhancement
@@
Table II and Fig. 5 illustrate that MKoIE significantly outperforms other methods in LLIE. Specifically, MKoIE demonstrates superior performance in image clarity and detail restoration, producing images with higher fidelity in terms of structural integrity while effectively reducing blurring and artifacts. LIME enhances visibility with noticeable improvements in color contrast and haze removal but introduces slight color distortions. SDD, while improving saturation in building areas and achieving a warmer appearance, suffers from blur and detail loss, particularly at the edges of buildings and within crowded regions. Both SMNet and EFINet exhibit over-sharpening effects, causing certain image details to appear excessively pronounced. In contrast, MKoIE generates images that closely resemble real-world scenes, with textures and colors more consistently preserved, thereby achieving a balance between enhancement and realism.
V-B3 Nighttime Haze Enhancement
@@
The evaluation of MKoIE was conducted using the RESIDE and CDD datasets, consistent with those in daytime ID and LLIE tasks. As shown in Fig. 5, while IENHC and ECNHI partially remove haze, residual haze remains visible, particularly in the background regions. In contrast, MKoIE cexhibits exceptional performance in color restoration, producing images with a natural tone devoid of oversaturation or color distortion, thereby enhancing both visual comfort and scene realism. In addition, MKoIE demonstrates superior detail preservation, with ground textures in the foreground and building outlines in the background clearly discernible. Quantitative evaluations in Table III further substantiates MKoIE’s superiority, as it achieves the best results in all metrics, underscoring its outstanding recovery performance under low-light and hazy conditions. Fig. 6’s histogram shows that our method achieves a balanced intensity distribution, along with superior brightness, color, and contrast enhancements. MKoIE produces a broad, well-spread histogram, indicating an optimal use of the intensity range. This corresponds to better brightness, more natural color distribution, and higher contrast compared to the other methods. Similarly, we conducted a detail enhancement comparison experiment, as shown in Fig. 7. The original image contains significant haze, especially around the object. IENHC yields a moderate clarity improvement: text becomes more legible, but background details remain largely unrecovered. MRP and OFSD both reduce haze and enhance image contrast and color saturation, though their results exhibit a slight remaining color cast. CEEF achieves only modest haze removal with a mild increase in contrast. In contrast, MKoIE shows pronounced improvements in overall color, contrast, and brightness. It also excels at restoring fine details, as evidenced by the clear textures visible in the red-boxed zoomed region in Figure 7. These visual enhancements correspond with MKoIE’s top quantitative metrics in Table III, confirming the method’s superior performance across both objective and subjective evaluations.
V-C Real-world Degradation Analysis
@@
Figure 8: Visual comparisons on real-world degradation scenes. Zooming-in is recommended for a better visualization and comparisons.
@@
TABLE IV: Ablation Studies on the TNL in ID, LLIE and NHIE tasks.
@@
TABLE V: Ablation Studies on the MRFE in ID, LLIE and NHIE tasks.
@@
TABLE VI: Ablation Studies on the Different Loss Function in ID, LLIE and NHIE tasks.
@@ V-D Ablation Study
V-D1 Ablation Studies on TNL
To validate the effectiveness of the proposed TNL modules, we assign different image restoration tasks to NHIE sub-node to train another version of MKoIE, as shown in Fig. 9, only a single sub-network was used to complete the three tasks. As shown in Table IV, compared to the full version of MKoIE, the performance of ID, LLIE, and NHIE degrades when the TNL module is specified independently. This highlights the effectiveness of the proposed shared and task-specific feature extraction methods. By leveraging its distinct architecture, the TNL module is designed to enhance feature extraction and thus improve overall task performance. It integrates the fundamental features required for ID, LLIE, and NHIE, without TNL module, MKoIE exhibits poor performance in the three tasks.
V-D2 Ablation Studies on Network Modules
To assess the significance of the MRFE module in MKoIE, we performed an ablation study by completely removing it. As shown in Table V, its removal consistently led to declines in PSNR and SSIM across all tasks. This demonstrates that the absence of the MRFE module limits the network’s ability to extract essential features, resulting in reduced restoration performance.
V-D3 Ablation Studies on Loss Function
Each component of the hybrid loss function show in subsection IV-F. To demonstrate the effectiveness of hybrid loss function, we train two versions: one using only the reconstruction loss and the other using only the perceptual loss. From Table VI, the declines in both PSNR and SSIM indicate that hybrid loss function guide the network toward improved restoration performance. Reconstruction and perceptual loss enhances restoration performance by balancing numerical stability, outlier resilience, perceptual consistency, and semantic preservation, achieving optimal results.
V-E Improving SOD Tasks with MKoIE
TABLE VII: The comparison for SOD task between original images and images enhanced by the MKoIE includes MAE, max F-measure($F_{\beta}^{\text{max}}$), max E-measure($E_{\phi}^{\text{max}}$)and S-measure($S_{\alpha}$).
@@ V-E Improving SOD Tasks with MKoIE
We utilized U2Net [48] and show its performance on CDD dataset. Notably, we directly applied the pretrained U2Net model without further training, achieving a marked enhancement in SOD through MKoIE preprocessing. As shown in Fig. 10 and Table VII, the U2Net model achieved significant improvements in MKoIE-processed images, enabling more accurate detection and segmentation of salient objects. This result validates the effectiveness of MKoIE, demonstrating its practical application value in traffic scenarios under complex weather conditions. Specifically, in autonomous driving and traffic monitoring systems, it enhances the detection and segmentation performance of critical targets such as pedestrians and vehicles, thereby further improving traffic safety.
V-F Complexity and Running Time Comparison
TABLE VIII: We selected the top four models based on the quantitative comparison of the NHIE task and present the corresponding average runtime (in seconds).
@@ V-F Complexity and Running Time Comparison
MKoIE adopts an efficient structural design, which can significantly improve the image recovery speed in low light and haze environments. We randomly selected 50 images from the test set and conducted tests under a consistent hardware setup. As shown in Table VIII, the average processing time of our model is about 1.68 seconds. The results show that MKoIE has advantages in both efficiency and effectiveness, demonstrating its potential for deployment in time-sensitive ITS scenarios such as traffic monitoring and autonomous driving.
VI Conclusions
This paper proposes a multi-knowledge-oriented imaging enhancer (MKoIE) to address salient object detection challenges in ITS under adverse imaging conditions. MKoIE effectively handles three types of degradation (i.e., daytime haze, low light, and nighttime haze) through a task-oriented node learning mechanism and self-attention module, while achieving efficient multi-scale feature extraction through a multi-receptive field enhancement module, making the method highly suitable for real-time ITS applications. Experimental results demonstrate that the proposed MKoIE achieves superior performance across various weather and imaging conditions, thus providing an effective solution for image enhancement and salient object detection in challenging environments, contributing to safer and more reliable transportation systems.